# Insights from the Usage of the Ansible Lightspeed Code Completion Service

### v2

@@ Abstract.
The availability of Large Language Models (LLMs) which can generate code, has made it possible to create tools that improve developer productivity. Integrated development environments or IDEs which developers use to write software are often used as an interface to interact with LLMs. Although many such tools have been released, almost all of them focus on general-purpose programming languages. Domain-specific languages, such as those crucial for Information Technology (IT) automation, have not received much attention. Ansible is one such YAML-based IT automation-specific language. Ansible Lightspeed is an LLM-based service designed explicitly to generate Ansible YAML given natural language prompt.
This paper first presents the design and implementation of the Ansible Lightspeed service. We then evaluate its utility to developers using diverse indicators, including extended utilization, analysis of user rejected suggestions, as well as analysis of user sentiments. The analysis is based on data collected for 10,696 real users including 3,910 returning users. The code for Ansible Lightspeed service and the analysis framework is made available for others to use.
To our knowledge, our study is the first to involve thousands of users in evaluating code assistants for domain-specific languages. We propose an improved version of user acceptance rate and we are the first code completion tool to present N-Day user retention figures. With our findings we provide insights into the effectiveness of small, dedicated models in a domain-specific context. We hope this work serves as a reference for software engineering and machine learning researchers exploring code completion services for domain-specific languages in particular and programming languages in general.
Large Language Models, Generative Models, Code Completion, IDE, User Study, Ansible
$c$s: Computing methodologies Natural language processingccs: Software and its engineering Context specific languagesccs: Software and its engineering Application specific development environments
@@ 1. Introduction
Ansible (Ansible Community, 2024) is a YAML based domain-specific language dedicated to IT automation. It is one of the most widely used infrastructure as code tools (Guerriero et al., 2019; StackExchange, 2023). Since it is open-source, we do not know the precise number of developers who use Ansible but we can get an indication of its popularity with the Ansible python community package (Team, 2024b) getting over 5 million downloads per month (Flynn, 2024). Ansible also has an active community on GitHub (Team, 2024a) with about 5,000 contributors and about 54,000 commits on a yearly basis, which is likely a small fraction of the actual user base. [^1] Another indication of Ansible’s popularity is from Ansible’s social media accounts of Reddit (Team, 2024c) and X (Team, 2024d), counting 63,000 readers and 72,000 followers, respectively.
@@
Code generation models, and some of the systems that utilize them, have emerged as powerful tools for software developers and system analysts (GitHub, 2021; Tab9, 2018; Replit, 2016; Amazon, 2023). Studies show that AI tools for general-purpose programming languages improve productivity and also suggest that novice programmers may benefit more from such tools (Weber et al., 2024; Peng et al., 2023; Barke et al., 2023). However, generic models do not perform as well on domain-specific languages. Pujar et al. (2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on benchmark Natural Language to Ansible-task generation test data. But this work does not involve real users in their development environment interacting with a model.
@@
Ansible Lightspeed was available since April 15 2023 to a limited number of users as part of a closed beta release, and then generally to everyone as part of a free Tech Preview version from June 2023  (RedHat, 2023) until May 15 2024. During this period thousands of developers interacted with the system.
In this paper we analyze the interactions and feedback of 10,696 users of the closed beta and free Tech Preview versions of Ansible Lightspeed, who consented to share their data. To our knowledge, there has been no published study, with several thousand users, on the use of code assistants for domain-specific languages.
@@
  • We propose an improved acceptance criteria that checks if the accepted suggestion was actually used by the user.
@@ 2.2. Inference Pipeline
WCA-Ansible is a transformer-based decoder trained from scratch on natural language, source code, and Ansible data. We use the Hugging Face Transformers framework (Wolf et al., 2020) to train a 350 million parameters model with a context window of size 1024; this model utilizes fewer parameters than the model available to subscription holders. The model is pre-trained with about 143 billion tokens and then fine-tuned with about 500k natural language to Ansible task samples. The fine-tuning data is created using a technique similar to the one of Ansible Wisdom (Pujar et al., 2023). The pre-training and fine-tuning data is license filtered to remove all data samples coming from files or repositories with restrictive licenses. We also filter the data to remove hate, abuse and profanity. The pre-training data is also deduplicated at a file level. The model server runs on machines containing A100100100100 GPUs equipped with sufficient memory to return an inference response within $22$ seconds on average.
@@
After the inference response is received by the Ansible Lightspeed service the response is again passed through the Ansible Anonymizer to cleanup any personal identifiable information that might be present in the model output. To ensure the suggestion follows the good practices of Ansible content development defined by the Ansible community (Community, 2019), the response is passed through Ansible Risk Insights or ARI tool (et al, 2020). The ARI tool runs multiple custom rules to modify the suggestion as required to ensure best practices such as invoking modules using their fully-qualified collection names and reusing Ansible variables defined in the context. It also redacts any credentials and removes and/or replaces any deprecated Ansible syntax from the suggestions. After this post-processing stage is complete, the response is sent back to the client.
@@ 2.4. Data Collection and Analysis Framework
An event tracking platform monitors and records a wide range of events about user interactions, and system activities. This platform is a critical part of our data collection pipeline, ensuring that we systematically capture important events within our system. It includes user-initiated actions like completion requests, suggestion acceptance, as well as service-related events like server requests, matching sources, and post-processing logs.
@@ 3.1. Temporal Analysis
Lightspeed’s goal is to improve the productivity of developers who use Ansible for work. Understanding when requests were made could give some hint that Lightspeed is being used for work related queries.
@@
In Figure 6 we can also see a significant decline in the completion requests between December 23, 2023, and January 12, 2024. This could be due to the Holiday season when many offices are usually closed.
@@ 3.3. Edit Analysis
In 24,811 instances (60.60% of the accepted suggestions), Ansible Lightspeed’s suggestions were used without edits so we can assume that they aligned with user requirements. In 5,672 scenarios (13.85% of the accepted suggestions), users made small adjustments, less than 50% edits [^6] Suggestion analyzer part of Sec. 2.4, which explains how edits are calculated., to the accepted suggestions. An example of such a light or minor edit would be, when Ansible Lightspeed accurately predicts the correct module, users modified one or more module keys (options) and/or values to better align with their specific needs.
@@
We also noticed that users deleted 7,436 suggestions after acceptance, which is 18.16% of the total initially accepted suggestions. Currently there is no way to tell why the users would delete accepted suggestions and move on to next request. One explanation is that users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to initially accept and then delete after closer review. Another possible reason could be that they were experimenting with different prompts trying one after another and discarding the answers.
@@ 3.4. User Acceptance Rate
The edit analysis described above shows that the acceptance rate as currently defined can be misleading. Which is why we define a new acceptance rate metric, called Edit¡50% acceptance rate. Edit¡50% Acceptance rate is the percentage of accepted suggestions, in which we consider a suggestion accepted only if the following conditions are fulfilled:
@@
As per this definition, the Edit¡50% acceptance rate for Ansible Lightspeed comes out to 49.08%, compared to the initial acceptance rate of 65.92%. We feel Edit¡50% acceptance rate more closely aligns with users perception of what they consider accepted, compared to the initial acceptance rate.
@@
Ansible Lightspeed generates multi-line Ansible code suggestions. Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and Edit¡50% acceptance rate.
Table 2. Comparison of acceptance rate of Ansible Lightspeed with the other code completion tools. Google MLECC (Maxim Tabachnyk, 2022) stands for ML enabled code completion. Ansible Lightspeed Initial accept are percent suggestions which were accepted by users. Edit¡50% are percent suggestions which were accepted and had less than 50% user edits, based on difflib SequenceMatcher(Library, 2001).
@@
Edit¡50% (Ansible) 49.1% 3,910
Q: At what rate do users accept suggestions? Initial acceptance rate is misleading since users can delete or heavily modify suggestions. 65.9% of suggestions are initially accepted. We define new acceptance rate, Edit¡50% Acceptance rate. 49.1% of model suggestions were accepted as per Edit¡50% Acceptance rate. Ansible Lightspeed has relatively higher acceptance for Ansible, compared to more general tools for various programming languages.
3.5. Rejected Suggestion Analysis
There are a total of 1,710 initially accepted suggestions which were explicitly edited by the users at the module level. We consider such suggestions to be rejected for Edit¡50% Acceptance rate. However, the users must have found something useful in the suggestion for them to accept it and then spend sometime editing it. We analyse these edits in order to better understand how the model suggestions differ from user expectations.
3.5.1. User preference
@@ 3.5.2. YAML reorganization
3.5.3. Functionally similar modules
@@
There are another 586 (28%) instances where the user edited the generated module. Understanding these edits requires further analysis.
@@ 4. User Feedback
A total of 605 users have provided rating and comments using the the sidebar for Ansible Lightspeed in the VS Code Editor, as shown in Figure 4.
@@ 4.2. User Comments
Apart from star-rating, users also submit feedback in the form of textual comments. By analyzing these comments we hope to find areas for improvement and also know what users like about the service. We manually review and classify negative feedback (1 or 2 stars) and positive feedback (4 or 5 stars) into various reasons. We also share some examples of real user comments, both positive and negative. These quotes have been rephrased for clarity and anonymity, removing any references to user details.
@@
“Has really bad recommendations. Tried creating a playbook with name ”Get …”” and it get me a module recommendation”
Finally, 5.79% of users are dissatisfied with their overall experience, specifically citing the need to re-authenticate periodically as an issue. The remaining 12.04% of comments are not informative.
@@
Breakdown of reasons for positive experience is presented in Figure 11. Most users, about 42.7%, consider improved productivity as Ansible Lightspeed’s key benefit. Multiple users mentioned that they noticeably saved time by using Lightspeed.
@@ 7. Conclusion and Future work
We hope that our proposed user edit-based acceptance criteria are used by other code completion tools to get a better estimate of whether users truly found the model suggestion helpful. We also hope that by sharing Ansible Lightspeed’s user retention figures, others get encouraged to share similar figures for their code completion tools. And finally, by sharing the code for Ansible lightspeed service [^7] https://github.com/ansible/ansible-ai-connect-service and our analysis framework [^8] https://github.com/ansible-community/lightspeed-analysis-framework we hope to help others trying to develop similar tools.

### v4

@@ Abstract.
The availability of Large Language Models (LLMs) which can generate code, has made it possible to create tools that improve developer productivity. Integrated development environments or IDEs which developers use to write software are often used as an interface to interact with LLMs. Although many such tools have been released, almost all of them focus on general-purpose programming languages. Domain-specific languages, such as those crucial for Information Technology (IT) automation, have not received much attention. Ansible is one such YAML-based IT automation-specific language. Ansible Lightspeed is an LLM-based service designed explicitly to generate Ansible YAML, given natural language prompt.
In this paper, we present the design and implementation of the Ansible Lightspeed service. We then evaluate its utility to developers using diverse indicators, including extended utilization, analysis of user edited suggestions, as well as user sentiments analysis. The evaluation is based on data collected for 10,696 real users including 3,910 returning users. The code for Ansible Lightspeed service and the analysis framework is made available for others to use.
[…]
Large Language Models, Generative Models, Code Completion, IDE, User Study, Ansible, Acceptance Rate, IT Automation
@@ 1. Introduction
Ansible (Ansible Community, 2024) is a YAML based domain-specific language dedicated to IT automation. It is one of the most widely used infrastructure as code tools (Guerriero et al., 2019; StackExchange, 2023). Since it is open-source, we do not know the precise number of developers who use Ansible but we can get an indication of its popularity with the Ansible python community package (Team, 2024b) getting over 5 million downloads per month (Flynn, 2024). Ansible also has an active community on GitHub (Team, 2024a) with about 5,000 contributors and about 54,000 commits on an yearly basis, which is likely a small fraction of the actual user base. [^1] Another indication of Ansible’s popularity is from Ansible’s social media accounts of Reddit (Team, 2024c) and X (Team, 2024d), counting 63,000 readers and 72,000 followers, respectively.
@@
Code generation models, and some of the systems that utilize them, have emerged as powerful tools for software developers and system analysts (GitHub, 2021; Tab9, 2018; Replit, 2016; Amazon, 2023). Studies show that AI tools for general-purpose programming languages improve productivity and also suggest that novice programmers may benefit more from such tools (Weber et al., 2024; Peng et al., 2023; Barke et al., 2023). However, generic models do not perform as well on domain-specific languages. Pujar et al. (2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on benchmark Natural Language to Ansible-task generation test data. But this work does not involve real users in their development environment.
@@
Ansible Lightspeed was available since April 15 2023 to a limited number of users as part of a closed beta release, and then generally to everyone as part of a free Tech Preview version from June 2023  (RedHat, 2023) until May 15 2024. During this period thousands of developers interacted with the system. In this paper we analyze the interactions and feedback of 10,696 users of the closed beta and free Tech Preview versions of Ansible Lightspeed, who consented to share their data. To our knowledge, there has been no published study, with several thousand users, on the use of code assistants for domain-specific languages.
@@
  • We propose an improved acceptance criteria, called Strong Acceptance rate, that checks if the accepted suggestion was actually used by the user.
@@ 2.2. Inference Pipeline
WCA-Ansible is a transformer-based decoder trained from scratch on natural language, source code, and Ansible data. We use the Hugging Face Transformers framework (Wolf et al., 2020) to train a 350 million parameters model with a context window of size 1024; this model utilizes fewer parameters than the model available to subscription holders. The model is pre-trained with about 143 billion tokens and then fine-tuned with about 500k natural language to Ansible task samples. The training environment, hyper-paramters and fine-tuning data creation techniques are similar to Ansible Wisdom (Pujar et al., 2023). The pre-training and fine-tuning data is license filtered to remove all data samples coming from files or repositories with restrictive licenses. We also filter the data to remove hate, abuse and profanity. The pre-training data is also deduplicated at a file level. The model server runs on machines containing A100100100100 GPUs equipped with sufficient memory to restrict the end-to-end latency within $22$ seconds on average. Every model suggestion is a multi-line Ansible task with 6.13 lines and 20.1 tokens on average.
@@
After the inference response is received by the Ansible Lightspeed service the response is again passed through the Ansible Anonymizer to cleanup any personal identifiable information that might be present in the model output. To ensure the suggestion follows the good practices of Ansible content development defined by the Ansible community (Community, 2019), the response is passed through Ansible Risk Insights or ARI tool (et al, 2020). The ARI tool runs multiple custom rules to modify the suggestion as required to ensure best practices such as invoking modules using their fully-qualified collection names and reusing Ansible variables defined in the context. It also redacts any credentials and removes and/or replaces any deprecated Ansible syntax from the suggestions. After post-processing is complete, the response is sent back to the client.
@@ 2.4. Data Collection and Analysis Framework
The event tracking platform monitors and records a wide range of events about user interactions, and system activities. This platform is a critical part of our data collection pipeline, ensuring that we systematically capture important events within our system. It includes user-initiated actions like completion requests, suggestion acceptance, as well as service-related events like server requests, matching sources, and post-processing logs.
@@ 3.1. Temporal Analysis
Lightspeed’s goal is to improve the productivity of developers who use Ansible for work. Understanding when requests were made could give hint that Lightspeed is being used for work.
@@
In Figure 6 we can also see a significant decline in the completion requests between December 23, 2023, and January 12, 2024. This could be due to the Holiday season when many offices are closed.
@@ 3.3. Edit Analysis
Average number of lines per suggestion 6.13
Average number of tokens per suggestion 20.1
@@
In 24,811 instances (60.60% of the accepted suggestions), Ansible Lightspeed’s suggestions were used without edits so we can assume that they aligned with user requirements. In 5,672 scenarios (13.85% of the accepted suggestions), users made small adjustments, less than 50% edits, [^6] Suggestion analyzer part of Sec. 2.4 explains how edits are calculated. to the accepted suggestions. An example of such a light or minor edit would be, when Ansible Lightspeed accurately predicts the correct module, users modified one or more module keys (options) and/or values to better align with their specific needs.
@@
We also noticed that users deleted 7,436 suggestions after acceptance, which is 18.16% of the total initially accepted suggestions. Currently there is no way to tell why the users would delete accepted suggestions and move on to next request. One explanation is that users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to initially accept and then delete after closer review. Another reason could be that they were experimenting with different prompts trying one after another and discarding the answers.
@@ 3.4. User Acceptance Rate
The edit analysis described above shows that the acceptance rate as currently defined can be misleading. Which is why we define a new acceptance rate metric, called Strong Acceptance rate. Strong Acceptance rate is the percentage of accepted suggestions, where a suggestion is considered accepted only if:
@@
As per this definition, the Strong Acceptance rate for Ansible Lightspeed comes out to 49.08%, compared to the initial acceptance rate of 65.92%. We feel Strong Acceptance rate more closely aligns with users perception of what they consider accepted, compared to the initial acceptance rate.
@@
Ansible Lightspeed always generates multi-line Ansible code suggestions, with an average of 6.13 lines and 20.1 tokens per suggestion over all the suggestions shown in Table. 1. Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and Strong Acceptance rate.
Table 2. Comparison of acceptance rate of Ansible Lightspeed with the other code completion tools. Google MLECC (Maxim Tabachnyk, 2022) stands for ML enabled code completion. Ansible Lightspeed Initial accept are percent suggestions which were accepted by users. Strong Acceptance rate is percent suggestions which were accepted and had less than 50% user edits, based on difflib SequenceMatcher(Library, 2001).
@@
Strong accept (Ansible) 49.1% 3,910
Q: At what rate do users accept suggestions? Initial acceptance rate is misleading since users can delete or heavily modify suggestions. 65.9% of suggestions are initially accepted. We define new acceptance rate, Strong Acceptance rate. 49.1% of model suggestions were accepted as per Strong Acceptance rate. Ansible Lightspeed has relatively higher acceptance for Ansible, compared to more general tools for various programming languages.
3.5. Edited Module Analysis
There are a total of 1,710 initially accepted suggestions which were explicitly edited by the users at the module level. We consider such suggestions to be rejected for Strong Acceptance rate. However, the users must have found something useful in the suggestion for them to accept it and then spend sometime editing it.
We analyse these edits in order to better understand how the model suggestions differ from user expectations. The edit categories were identified based on manual analysis. The number of samples in each category was obtained based on rules to find similar samples. A suggestion could have been edited for multiple reasons, so an edited sample can belong to multiple categories.
User preference
@@ 3.5. Edited Module Analysis
Functionally similar modules
@@
There are another 586 (34.2%) cases where the user edited the module. Understanding these edits requires more analysis.
@@ 4. User Feedback
A total of 605 users have provided rating and comments using the the sidebar for Ansible Lightspeed in the VS Code Editor (Figure 4).
@@ 4.2. User Comments
Apart from star-rating, users also submit feedback in the form of textual comments. We manually review and classify negative feedback (1 or 2 stars) and positive feedback (4 or 5 stars). The categories are discovered in the data using open coding. Each comment is assigned to only one category, the main point of the feedback. We also share some examples of real user comments, both positive and negative. These quotes have been rephrased for clarity and anonymity, removing any references to user details.
@@
“Has really bad recommendations. Tried creating a playbook with name ”Get …” and it get me a module recommendation”
Finally, 5.79% of users are dissatisfied with their overall experience, specifically citing the need to re-authenticate periodically as an issue.
The remaining 12.04% of comments are not informative.
@@
Breakdown of reasons for positive experience is presented in Figure 11. About 42.7% of users consider improved productivity as Ansible Lightspeed’s key benefit. Multiple users mentioned that they noticeably saved time by using Lightspeed.
@@ 7. Conclusion and Future work
We hope that our proposed user edit-based acceptance criteria, Strong Acceptance rate, is used by other code completion tools to get a better estimate of whether users truly found the model suggestion helpful. We also hope that by sharing Ansible Lightspeed’s user retention figures, others get encouraged to share similar figures for their code completion tools. And finally, by sharing the code for Ansible lightspeed service [^7] https://github.com/ansible/ansible-ai-connect-service and our analysis framework [^8] https://github.com/ansible-community/lightspeed-analysis-framework we hope to help others trying to develop similar tools.